write an equation of the line that passes through each pair of points in slope-intercept form
(0,5) and (2,6)

Answer by addingup(3677)

(0,5) and (2,6)
y = mx+b
y_2-y_1 = m(x_2-x_1)
(y_2-y_1)/(x_2-x_1) = m flip the equation so we have the unknown on the right:
m = (y_2-y_1)/(x_2-x_1)
m = (6-5)/2-0)
m = 1/2 this is our slope. Now let's use the point slope formula with one of our points, any one of the two. I'll use (0,5):
y-5 = 1/2(x-0) multiply times 2
2y-10 = x-0
2y = x-0+10
2y = x+10
---------------------
Just to amuse you, I'll do it with the other set of points (2,6):
y-6 = 1/2(x-2)
2y-12 = x-2
2y = x-2+12
2y = x+10 As you see, the result is exactly the same
:
John